# Emission and Absorption Lines in Photospheric Radius Expansion Bursts of 4U 1820–30

### v1

@@
Emission and Absorption Lines in Photospheric Radius Expansion Bursts of 4U 1820–30
@@
We analyze the emission and absorption lines during photospheric radius expansion (PRE) X-ray bursts from the ultracompact binary 4U 1820–30, observed with the Neutron Star Interior Composition Explorer (NICER ). Using Monte Carlo simulations to estimate the significance, we identified a 1 keV emission line from 14 bursts, a 3 keV absorption line from 12 bursts, and 1.6 keV absorption from one burst. By coadding the burst spectra at the maximum radius phase, we detected a 1.034 keV emission line with significance of $14.2\sigma$, and absorption lines at 1.64 and 3 keV with significances of $10.8\sigma$ and $11.7\sigma$, respectively. The observed energy shifts are consistent with the prediction from the burst-driven wind model, indicating that all three spectral features are produced by the PRE wind. Analysis of the ratios between the emission and absorption line energies suggests that the 1 keV feature is a superposition of several narrower Fe L-shell lines. To evaluate the scientific capabilities of the Hot Universe Baryon Surveyor (HUBS), we simulated mock observations of multiple narrow lines near 1 keV. The results demonstrate that HUBS is well suited for detailed studies of the 1 keV emission line during bursts, offering significant potential to advance our understanding of these phenomena.
@@ 1 Introduction
4U 1820–30 is a persistent ultracompact X-ray binary (UCXB) with an orbital period of 11.4 minutes (Stella et al., 1987). Due to its short orbital period, the donor star is likely a white dwarf with a mass of $\sim 0.06-0.07\leavevmode\nobreak\ M_{\odot}$ (Paczynski & Sienkiewicz, 1981; Nelson et al., 1986; Rappaport et al., 1987). In this system, a neutron star (NS) accretes helium-rich material from its companion, potentially triggering unstable thermonuclear burning, resulting in type I X-ray bursts (or simply X-ray burst; Lewin et al., 1993; Strohmayer & Bildsten, 2006; Falanga et al., 2008; Galloway et al., 2008; Galloway & Keek, 2021). These bursts typically release a total energy of $\sim 10^{39}$ erg over 10–100 s. Occasionally, bursts can exceed the Eddington limit, causing the NS surface layers to lift and resulting in photospheric radius expansion (PRE; Lewin et al., 1993). In rare cases, the photosphere radius may expand by a factor of $>100$, which is denoted as “superexpansion” burst (in’t Zand & Weinberg, 2010; Yu et al., 2024).
@@
A strong PRE burst can drive the matter outflow from the NS surface, such as generating a super-Eddington wind (Paczynski & Proszynski, 1986; Yu & Weinberg, 2018; Guichandut et al., 2021). There is some evidence suggesting that the wind is polluted with heavy nuclear burning ashes (in’t Zand & Weinberg, 2010; Yu & Weinberg, 2018; Guichandut et al., 2021), as type I X-ray bursts are known sites for the production of heavy element via the $rp$-process (Schatz et al., 2003). Convection at the onset of the burst can mix heavy elements from the deeper burning layers, which are then ejected by the wind and exposed at the photosphere (Weinberg et al., 2006; Yu & Weinberg, 2018). In superexpansion bursts, the column density of ejected ashes is expected to be particularly high due to the energetic nature of these events. As a result, the heavy element ashes may be observed as spectral lines or absorption edges (in’t Zand & Weinberg, 2010). For example, in two superexpansion bursts from 4U 0614+091 and 4U 1722–30 detected by RXTE, in’t Zand & Weinberg (2010) reported spectral edges in the range of 6–11 keV. These edge energies can be associated with a redshifted hydrogen/helium-like Ni edge, suggesting the presence of heavy-element ashes in the wind. Li et al. (2018) identified an absorption edge in a PRE burst from GRS 1747–312 during RXTE observations . The edge energy, remaining at $\approx 8\leavevmode\nobreak\ {\rm keV}$ during the cooling tail, was attributed to photoionization absorption edges of hydrogen-like nickel, redshifted to match the observed energy. Based on this, they estimate the NS mass, radius and surface gravitational redshift factor.
@@
In 2017, Keek et al. (2018) reported a strong PRE burst in 4U 1820–30 from NICER observations. In addition to a blackbody component, an extra optically thick Comptonization model has to be added to fit the time-resolved burst spectra. Four of these bursts exhibited significant spectral features, including an emission line at 1 keV and absorption features at 1.7 keV and 3 keV (Strohmayer et al., 2019). They identified a relative spectral shift of approximately 1.046 in bursts with a photospheric radius of around 75 km, attributed to the combined effects of gravitational redshift and Doppler blueshift in a burst-driven wind. However, from the calculations in Guichandut et al. (2021), only a static envelope can be produced for such a small photospheric radius, and the wind-induced blueshift is negligible. More recently, Yu et al. (2024) analyzed 15 type I X-ray bursts from 4U 1820–30 using NICER (see also Jaisawal et al., 2024). The time-resolved burst spectra revealed that all these bursts were PRE with photospheric radii ranging from 55 to $10^{3}$ km. Additionally, they provided direct evidence that the accretion disk was distorted by the burst radiation by fitting the burst spectra with a reflection model.
@@
In this work, beyond the analysis of Strohmayer et al. (2019) and Yu et al. (2024), we focus on the spectral lines of 15 X-ray bursts from 4U 1820–30 observed by NICER. This work is organized as follows. In Section 2, we perform the spectral lines fitting during the X-ray bursts. We discuss and summarize the results in Sections 3 and 4, respectively.
@@ 2 Observations and data reduction
NICER observed 4U 1820–30 from 2017 to present for a total exposure time of 812 ks, in which 15 type I X-ray bursts were detected with a short duration and a high peak count rate. We processed all archived NICER data by applying the standard filtering criteria using HEASOFT V6.33.2 and the NICER Data Analysis Software (NICERDAS) V2.0.7. The 0.1 s binned light curves were extracted using nicerl3-lc from the calibrated unfiltered (UFA) files and cleaned events files, see Fig. 1. We extracted the spectra, ancillary response files (ARFs), response matrix files (RMFs) and the 3c50 background spectra (Remillard et al., 2021) using nicerl3-spect. We performed the spectra analysis using Xspec v12.14.1 (Arnaud, 1996). The errors of all parameters are quoted at the $1\sigma$ confidence level.
@@
All time-resolved burst spectra can be fitted well by using the enhanced persistent emission model ($f_{\rm a}$ model), adding a reflection component from the surrounding accretion disk (disk reflection model), or the double-blackbody model, as reported by Yu et al. (2024). All bursts showed the characteristics of PRE. However, the $f_{a}$ model failed to provide a physically motivated and self-consistent explanation for the observations. This study focuses on the spectral features of X-ray bursts from 4U 1820–30 as detected by NICER. Due to its phenomenological simplicity, the double-blackbody model was adopted to describe the continuum spectra for subsequent analysis. For more details on continuum fitting, see Yu et al. (2024).
Table 1: Summary of burst observations.
[…]
1 1050300108 2017-08-29 08:58:27 57994.37393 2.15 0.3-0.8
2 1050300108 2017-08-29 11:08:51 57994.46448 1.56 0.4-1.4
3 1050300109 2017-08-30 05:24:49 57995.22557 2.28 0.4-0.8
4 1050300109 2017-08-30 08:11:55 57995.34162 2.08 0.4-0.8
5 1050300109 2017-08-30 14:32:38 57995.60600 1.89 0.9-1.4
6 2050300104 2019-06-07 07:05:51 58641.29573 2.52 0.5-1.1
7 2050300108 2019-06-12 20:09:49 58646.84015 1.96 0.1-0.8
8 2050300110 2019-06-14 07:50:25 58648.32669 2.25 0.4-1.4
[…]
10 2050300119 2019-06-26 07:31:06 58660.31327 2.60 0.3-1.0
[…]
14 2050300124 2019-07-01 14:15:25 58665.59405 2.54 0.4-1.1
15 4680010101 2021-05-02 14:41:42 59336.61230 2.57 0.6-1.2
111
@@
Figure 1: The 0.1 s binned light curves of 15 X-ray bursts from 4U 1820–30 observed by NICER in 0.5–10 keV are marked as red lines. The right axis shows the 3–10 keV light curves as black lines. The persistent emissions are regarded as background and subtracted. The gray shaded region mark the intervals used to extract PRE phase spectra.
@@
Figure 2: The PRE phase spectra and residuals for burst # 9. Top panel: the PRE phase spectra from the burst # 9. The best-fit model, ${\tt Tbabs}\times({\tt bbodyrad+gaussion+bbodyrad})\times{\tt gabs}\times{\tt gabs}$, represents as a red solid line. The 1 keV emission line are clearly shown in the spectra, and a pair of absorption lines are indicated near 1.6 and 3 keV. The two middle panels are the residuals of the spectra fitted with 3c50 and SCORPEON backgrounds, respectively.
The bottom panel shows the residuals of the best-fit model to the spectra.
@@
Figure 3: The distributions of the $\Delta C$ of $10^{5}$ simulations for the lines of burst # $\#\ 9$. The distributions fitted by $\chi^{2}$-distribution with two dofs (red dashed-dotted lines). From top to bottom panel, the arrow marks the significance for the 1 keV emission line, 1.6 and 3.0 keV absorption lines, respectively.
@@ 2.1 Single burst spectra
Strohmayer et al. (2019) found that the spectral lines appeared only during the PRE phase. No spectral lines were present in the pre-burst persistent spectra and the initial and cooling phases of the bursts (Yu et al., 2024). Therefore, we extracted spectra during the PRE phase for each burst based on the time-resolved spectroscopy. The spectra were extracted with the exposure time varying from 0.3 to 2 s. The gray area in Fig. 1 marks the intervals used to extract PRE phase spectra. The burst spectra were grouped using grappha with a minimum count of 20 and regarded the persistent spectra as background and unchanged during bursts. The PRE spectra of most burst were dominated by the persistent emission above 5 keV, see the burst light curve in 5–10 keV after subtracting the persistent count rate in Fig. 1. To focus on searching for spectral features, we fitted each spectrum in 0.5–5 keV with the double blackbody model, i.e., TBabs $\times$ (bbodyrad + bbodyrad). The absorption column density was fixed as the best fitting value from the persistent spectra, $N_{\rm H}=1.93\times 10^{21}\leavevmode\nobreak\ {\rm cm}^{-2}$.
The double blackbody model provides a reasonable qualitative fit to the continuum spectra, but there are substantial spectral features evident in the residual. For each X-ray burst, we select a spectrum with significant emission and absorption lines to analyze the spectral lines of the burst, In Fig. 2, as an example, we show the PRE phase spectrum extracted from burst # 9, which has $\chi^{2}=315.82$ for 230 degrees of freedom (dof) fitted by the double blackbody model. The residuals showed apparent line features near 1.0 and 1.6 keV, and a possible 3 keV absorption line guided by Strohmayer et al. (2019). Adding one Gaussian (gauss) and two absorption components (gabs) to describe these features, i.e., the overall model is ${\tt Tbabs}\times({\tt bbodyrad+gaussion+bbodyrad})\times{\tt gabs}\times{\tt gabs}$, which can obviously improve the fitting result. The $\chi^{2}$ is 256.51 for 224 dof. The best fitted centroid energies and line width are $E_{1}=1.028\pm{0.024}$ keV and width $\sim 0.05$ keV, $E_{2}=1.625\pm{0.021}$ keV and width $\sim 0.02$ keV, and $E_{3}=2.894\pm{0.055}$ keV and width $\sim 0.04$, respectively.
To confirm the robustness of spectral features against background modeling uncertainties, we reprocessed the data using nicerl3-spec with the the SCORPEON background model  [^2] https://heasarc.gsfc.nasa.gov/docs/nicer/analysis__\__threads/scorpeon-xspec/. For burst # 9 as an example, we compared the spectral lines from different background models. The best-fitted parameters of the 1 keV emission and 3 keV absorption lines were consistent within $1\sigma$ confidence level compared with the 3c50 background model (see Fig. 2).
@@ 2.2 Spectra line significance assessment
The significance of spectral lines is not reliably determined using goodness-of-fit or F-test methods, particularly for emission and absorption lines (Protassov et al., 2002; Wang et al., 2022). The F-test tends to overestimate the significance of these features (Protassov et al., 2002; Lanzuisi et al., 2024). Therefore, we employed the Monte Carlo (MC) method to evaluate the significance of potential spectral lines (Porquet et al., 2004; Tombesi et al., 2010; Gofford et al., 2013). This approach has proven robust and has become the standard practice in the past decade (Li et al., 2022; Xu & Wang, 2022; Parra et al., 2023; Zhang et al., 2024)
@@
where $L_{\rm null}$ and $L_{\rm line}$ represent the likelihoods for the null hypothesis model and the model with an additional line component, respectively. The $\Delta C_{\rm true}$ for three spectral features are calculated as, $\Delta C_{{\rm true},\leavevmode\nobreak\ E_{1}}=33.93$, $\Delta C_{{\rm true},\leavevmode\nobreak\ E_{2}}=27.01$, and $\Delta C_{{\rm true},\leavevmode\nobreak\ E_{3}}=6.20$, respectively. To assess these features, we generate $10^{5}$ distributions of best-fitted parameters within the uncertainties using simpars command in Xspace. Each set of model parameters was loaded from the generated distribution, and a spectrum was simulated using the fakeit command in Xspec, preserving all observational parameters (exposure, response files, background) of the initial spectrum. The simulated spectra were first fitted with the double-blackbody model to obtain the baseline $C_{\rm null}$ values. Then, an emission or absorption line component was added to the model, and the spectra were re-fitted to obtain $C_{\rm line}$. The resulting $\Delta C_{\rm{sim}}$ values were recorded for each simulated spectrum, forming a distribution for statistical analysis. This method provides a robust estimate of the significance of the spectral features.
For the $10^{5}$ simulations, the distributions of $\Delta C_{\text{sim}}$ were fitted to a $\chi^{2}$-distribution, and the $p$-values were inferred based on these fits. Fig. 3 illustrates the distributions and fitting results for the three spectral features of burst # 9. The results show that the simulated distribution is consistent with the $\chi^{2}$-distribution with two dofs. The two dofs of $\chi^{2}$ are derived from the two free parameters of gaussian or absorption given the fixed line width. The derived significances are, $\sigma_{\rm E1}=5.5\sigma$, $\sigma_{\rm E2}=4.8\sigma$, and $\sigma_{\rm E3}=2.0\sigma$.
The MC method was subsequently applied to all spectra to assess the significance of detected lines. Only spectral lines with a significance greater than $2\sigma\ (95\%)$, were considered detections and included in the further analyses (see Table 2). For burst # 2, no spectral lines were detected, consistent with the findings of Strohmayer et al. (2019). In total, we identified the 1 keV emission line in 14 bursts, the 3 keV absorption line in 12 bursts, and the 1.6 keV absorption line in one burst (# 9). The spectra and the best-fitted model are shown in Fig. 4. Table 2 summarizes the results of best-fit parameters of the emission and absorption lines. The strongest identified feature is the 1 keV emission line from burst # $\#\ 8$ with a significance of $10.8\sigma$. All spectral features exhibit narrow line widths of approximately 0.02–0.08 keV; thus, the line width was fixed during fitting.
@@
Figure 4: The unabsorbed spectra extracted during the PRE phase for all bursts in 0.3–5 keV. The red solid lines are the best-fitting model, including the continuum and spectral lines if necessary. The residuals of the best fit model to the data are plotted in bottom panels.

### v2

# Emission and absorption lines in the photospheric radius expansion bursts of 4U 1820–30

@@
Emission and absorption lines in the photospheric radius expansion bursts of 4U 1820–30
@@
We analyze the emission and absorption lines during photospheric radius expansion (PRE) X-ray bursts from the ultracompact binary 4U 1820–30, observed with the Neutron Star Interior Composition Explorer (NICER). Using Monte Carlo simulations to estimate the significance, we identified a 1 kiloelectron volt (keV) emission line from 14 bursts, a 3 keV absorption line from 12 bursts, and a 1.6 keV absorption line from one burst. By coadding the burst spectra at the maximum radius phase, we detected a 1.034 keV emission line with significance of $14.2\sigma$, along with absorption lines at 1.64 and 3 keV, with significances of $10.8\sigma$ and $11.7\sigma$, respectively. The observed energy shifts are consistent with the prediction from the burst-driven wind model, indicating that all three spectral features are produced by the PRE wind. An analysis of the ratios between the emission and absorption line energies suggests that the 1 keV feature is a superposition of several narrower Fe L-shell lines. To evaluate the scientific capabilities of the Hot Universe Baryon Surveyor (HUBS), we simulated mock observations of multiple narrow lines near 1 keV. These results demonstrate that HUBS is well suited for detailed studies of the 1 keV emission line during bursts, offering significant potential to advance our understanding of these phenomena.
@@ 1 Introduction
The 4U 1820–30 system is a persistent, ultracompact, X-ray binary (UCXB) with an orbital period of 11.4 minutes (Stella et al., 1987). Due to its short orbital period, the donor star is likely a white dwarf with a mass of $\sim 0.06-0.07\leavevmode\nobreak\ M_{\odot}$ (Paczynski & Sienkiewicz, 1981; Nelson et al., 1986; Rappaport et al., 1987). In this system, a neutron star (NS) accretes helium-rich material from its companion, potentially triggering unstable thermonuclear burning, resulting in type I X-ray bursts (or simply X-ray burst; Lewin et al., 1993; Strohmayer & Bildsten, 2006; Falanga et al., 2008; Galloway et al., 2008; Galloway & Keek, 2021). These bursts typically release a total energy of $\sim 10^{39}$ erg over 10–100 s. Occasionally, bursts can exceed the Eddington limit, causing the NS surface layers to lift, resulting in photospheric radius expansion (PRE; Lewin et al., 1993). In rare cases, the photosphere radius can expand by a factor of $>100$, described as a “superexpansion” burst (in’t Zand & Weinberg, 2010; Yu et al., 2024).
@@
A strong PRE burst can drive the matter outflow from the NS surface, such as generating a super-Eddington wind (Paczynski & Proszynski, 1986; Yu & Weinberg, 2018; Guichandut et al., 2021). There is some evidence suggesting that the wind is polluted with heavy nuclear burning ashes (in’t Zand & Weinberg, 2010; Yu & Weinberg, 2018; Guichandut et al., 2021), as type I X-ray bursts are known sites for the production of heavy element via the $rp$-process (Schatz et al., 2003). Convection at the onset of the burst can mix heavy elements from the deeper burning layers, which are then ejected by the wind and exposed at the photosphere (Weinberg et al., 2006; Yu & Weinberg, 2018). In superexpansion bursts, the column density of ejected ashes is expected to be particularly high due to the energetic nature of these events. As a result, the heavy element ashes could be observed as spectral lines or absorption edges (in’t Zand & Weinberg, 2010). For example, in two superexpansion bursts from 4U 0614+091 and 4U 1722–30 detected by RXTE, in’t Zand & Weinberg (2010) reported spectral edges in the range of 6–11 kiloelectron volts (keV). These edge energies can be associated with a redshifted hydrogen or helium-like Ni edge, suggesting the presence of heavy-element ashes in the wind. Li et al. (2018) identified an absorption edge in a PRE burst from GRS 1747–312 during RXTE observations. The edge energy, remaining at $\approx 8\leavevmode\nobreak\ {\rm keV}$ during the cooling tail, was attributed to photoionization absorption edges of hydrogen-like nickel, redshifted to match the observed energy. Based on this finding, they estimated the NS mass, radius, and surface gravitational redshift factor.
@@
In 2017, Keek et al. (2018) reported a strong PRE burst in 4U 1820–30 from NICER observations. In addition to a blackbody component, an extra optically thick Comptonization model has to be added to fit the time-resolved burst spectra. Four of these bursts exhibited significant spectral features, including an emission line at 1 keV and absorption features at 1.7 keV and 3 keV (Strohmayer et al., 2019). They identified a relative spectral shift of approximately 1.046 in bursts with a photospheric radius of around 75 km, attributed to the combined effects of gravitational redshift and Doppler blueshift in a burst-driven wind. However, from the calculations in Guichandut et al. (2021), only a static envelope can be produced for such a small photospheric radius and the wind-induced blueshift is negligible. More recently, Yu et al. (2024) analyzed 15 type I X-ray bursts from 4U 1820–30 using NICER (see also Jaisawal et al., 2024). The time-resolved burst spectra revealed that all these bursts were PRE with photospheric radii ranging from 55 to $10^{3}$ km. Additionally, they provided direct evidence that the accretion disk was distorted by the burst radiation by fitting the burst spectra with a reflection model.
@@
In this work, extending the analysis by Strohmayer et al. (2019) and Yu et al. (2024), we focus on the spectral lines of 15 X-ray bursts from 4U 1820–30 observed by NICER. This work is organized as follows. In Sect. 2, we describe how we performed the spectral line fitting during the X-ray bursts. We discuss and summarize the results in Sects. 3 and 4, respectively.
@@ 2 Observations and data reduction
NICER observed 4U 1820–30 from 2017 to present for a total exposure time of 812 ks, in which 15 type I X-ray bursts were detected with a short duration and a high peak count rate. We processed all archived NICER data by applying the standard filtering criteria using HEASOFT V6.33.2 and the NICER Data Analysis Software (NICERDAS) V2.0.7. The 0.1 s binned light curves were extracted using nicerl3-lc from the calibrated unfiltered (UFA) files and cleaned events files, see Fig. 1. We extracted the spectra, ancillary response files (ARFs), response matrix files (RMFs), and the 3c50 background spectra (Remillard et al., 2022) using nicerl3-spect. We performed the spectra analysis using Xspec v12.14.1 (Arnaud, 1996). The errors of all parameters are quoted at the $1\sigma$ confidence level.
@@
All time-resolved burst spectra can be fitted well by using the enhanced persistent emission model ($f_{\rm a}$ model), adding a reflection component from the surrounding accretion disk (disk reflection model), or using the double-blackbody model, as reported by Yu et al. (2024). All bursts showed the characteristics of PRE. However, the $f_{a}$ model failed to provide a physically motivated and self-consistent explanation for the observations. This study focuses on the spectral features of X-ray bursts from 4U 1820–30, as detected by NICER. Due to its phenomenological simplicity, the double-blackbody model was adopted to describe the continuum spectra for subsequent analysis. For more details on continuum fitting, we refer to Yu et al. (2024).
@@
Figure 1: 0.1 s binned light curves of 15 X-ray bursts from 4U 1820–30 observed by NICER in 0.5–10 keV are marked as red lines. The right axis shows the 3–10 keV light curves as black lines. The persistent emissions are regarded as background and subtracted. The gray shaded region mark the intervals used to extract PRE phase spectra.
@@
Figure 2: PRE phase spectra and residuals for burst # 9. Top panel: PRE phase spectra from the burst # 9. The best-fit model, ${\tt Tbabs}\times({\tt bbodyrad+gaussion+bbodyrad})\times{\tt gabs}\times{\tt gabs}$, is represented by the solid red line. The 1 keV emission line are clearly shown in the spectra and a pair of absorption lines are indicated near 1.6 and 3 keV. The two middle panels are the residuals of the spectra fitted with 3c50 and SCORPEON backgrounds, respectively. The bottom panel shows the residuals of the best-fit model to the spectra.
@@
Figure 3: Distributions of the $\Delta C$ of $10^{5}$ simulations for the lines of burst # $\#\ 9$. The distributions fitted by $\chi^{2}$-distribution with two dofs (red dashed-dotted lines). From top to bottom panel, the arrow marks the significance for the 1 keV emission line, 1.6 and 3.0 keV absorption lines, respectively.
@@ 2.1 Single burst spectra
Strohmayer et al. (2019) found that the spectral lines appeared only during the PRE phase. No spectral lines were present in the pre-burst persistent spectra and the initial and cooling phases of the bursts (Yu et al., 2024). Therefore, we extracted spectra during the PRE phase for each burst based on the time-resolved spectroscopy. The spectra were extracted with the exposure time varying from 0.3 to 2 s. The gray area in Fig. 1 marks the intervals used to extract PRE phase spectra. The burst spectra were grouped using grappha with a minimum count of 20 and regarded the persistent spectra as background and unchanged during bursts. The PRE spectra of most burst were dominated by the persistent emission above 5 keV, see the burst light curve in 5–10 keV after subtracting the persistent count rate in Fig. 1. To focus on searching for spectral features, we fitted each spectrum in 0.5–5 keV with the double-blackbody model, namely, TBabs $\times$ (bbodyrad + bbodyrad). The absorption column density was fixed as the best fitting value from the persistent spectra, $N_{\rm H}=1.93\times 10^{21}\leavevmode\nobreak\ {\rm cm}^{-2}$.
The double-blackbody model provides a reasonable qualitative fit to the continuum spectra, but there are substantial spectral features evident in the residual. For each X-ray burst, we selected a spectrum with significant emission and absorption lines to analyze the spectral lines of the burst. In Fig. 2, as an example, we show the PRE phase spectrum extracted from burst # 9, which has $\chi^{2}=315.82$ for 230 degrees of freedom (dofs) fitted by the double-blackbody model. The residuals showed apparent line features near 1.0 and 1.6 keV, and a possible 3 keV absorption line guided by Strohmayer et al. (2019). Adding one Gaussian (gauss) and two absorption components (gabs) to describe these features; namely, the overall model is ${\tt Tbabs}\times({\tt bbodyrad+gaussion+bbodyrad})\times{\tt gabs}\times{\tt gabs}$, which can obviously improve the fitting result. The $\chi^{2}$ is 256.51 for 224 dofs. The best-fit centroid energies and line width are $E_{1}=1.028\pm{0.024}$ keV and width $\sim 0.05$ keV, $E_{2}=1.625\pm{0.021}$ keV and width $\sim 0.02$ keV, and $E_{3}=2.894\pm{0.055}$ keV and width $\sim 0.04$, respectively.
To confirm the robustness of spectral features against background modeling uncertainties, we reprocessed the data using nicerl3-spec with the SCORPEON background model  [^1] https://heasarc.gsfc.nasa.gov/docs/nicer/analysis__\__threads/scorpeon-xspec/. Using burst # 9 as an example, we compared the spectral lines from different background models. The best-fit parameters of the 1 keV emission and 3 keV absorption lines were consistent within $1\sigma$ confidence level compared with the 3c50 background model (see Fig. 2).
@@ 2.2 Spectra line significance assessment
The significance of spectral lines was not reliably determined using goodness-of-fit or F-test methods, particularly for the emission and absorption lines (Protassov et al., 2002; Wang et al., 2022). The F-test tends to overestimate the significance of these features (Protassov et al., 2002; Lanzuisi et al., 2024). Therefore, we employed the Monte Carlo (MC) method to evaluate the significance of potential spectral lines (Porquet et al., 2004; Tombesi et al., 2010; Gofford et al., 2013). This approach has proven robust and has become the standard practice in the past decade (Li et al., 2022; Xu & Wang, 2022; Parra et al., 2024; Zhang et al., 2024)
@@
where $L_{\rm null}$ and $L_{\rm line}$ represent the likelihoods for the null hypothesis model and the model with an additional line component, respectively. The $\Delta C_{\rm true}$ for three spectral features are calculated as, $\Delta C_{{\rm true},\leavevmode\nobreak\ E_{1}}=33.93$, $\Delta C_{{\rm true},\leavevmode\nobreak\ E_{2}}=27.01$, and $\Delta C_{{\rm true},\leavevmode\nobreak\ E_{3}}=6.20$, respectively. To assess these features, we generate $10^{5}$ distributions of best-fitted parameters within the uncertainties using simpars command in Xspace. Each set of model parameters was loaded from the generated distribution, and a spectrum was simulated using the fakeit command in Xspec, preserving all observational parameters (exposure, response files, background) of the initial spectrum. The simulated spectra were first fitted with the double-blackbody model to obtain the baseline $C_{\rm null}$ values. Then, an emission or absorption line component was added to the model and the spectra were re-fitted to obtain $C_{\rm line}$. The resulting $\Delta C_{\rm{sim}}$ values were recorded for each simulated spectrum, forming a distribution for statistical analysis. This method provides a robust estimate of the significance of the spectral features.
For the $10^{5}$ simulations, the distributions of $\Delta C_{\text{sim}}$ were fitted to a $\chi^{2}$-distribution, and the $p$-values were inferred based on these fits. Figure 3 illustrates the distributions and fitting results for the three spectral features of burst # 9. The results show that the simulated distribution is consistent with the $\chi^{2}$-distribution with two dofs. The two dofs of $\chi^{2}$ are derived from the two free parameters of gaussian or absorption given the fixed line width. The derived significances are, $\sigma_{\rm E1}=5.5\sigma$, $\sigma_{\rm E2}=4.8\sigma$, and $\sigma_{\rm E3}=2.0\sigma$.
The MC method was subsequently applied to all spectra to assess the significance of detected lines. Only spectral lines with a significance greater than $2\sigma\ (95\%)$ were considered detections and included in the further analyses (see Table 1). For burst # 2, no spectral lines were detected, consistent with the findings of Strohmayer et al. (2019). In total, we identified the 1 keV emission line in 14 bursts, the 3 keV absorption line in 12 bursts, and the 1.6 keV absorption line in one burst (# 9). The spectra and the best-fitted model are shown in Fig. 4. Table 1 summarizes the results of best-fit parameters of the emission and absorption lines. The strongest identified feature is the 1 keV emission line from burst # $\#\ 8$ with a significance of $10.8\sigma$. All spectral features exhibit narrow line widths of approximately 0.02–0.08 keV; thus, the line width was fixed during fitting.
@@
Figure 4: Unabsorbed spectra extracted during the PRE phase for all bursts in 0.3–5 keV. The solid red lines indicate the best-fitting model, including the continuum and spectral lines if necessary. The residuals of the best fit model to the data are plotted in bottom panels.